# Code Clone Detection via an AlphaFold-Inspired Framework

### v1

@@ 3.1. From Protein to Code: Anology of Sequence
Second, both protein sequences and token sequences exhibit inner-sequence interaction, where the meaning of each element is context-dependent and determined by its neighboring elements within the sequence. In a protein sequence, the role of an amino acid is dependent on its neighboring residues. For instance, in the lysozyme protein sequence, the segment from positions $25$ to $35$ forms an $\alpha$-helix structure and consists of the following amino acids:
@@
where every letter denotes an amino acid and the accompanying number specifies its position. The relative positioning of amino acids such as W28 and W31 contributes to the formation of a hydrophobic core that stabilizes the $\alpha$-helix structure, while the local arrangement of W28–V29–A30 creates a hydrophobic patch that is essential for local stability and contributes to the protein’s overall function. Similarly, in a token sequence, the semantic role of each token is determined by its surrounding token context within the sequence, as program semantics often arise from token patterns rather than individual tokens. For example, considering the C‑style loop
@@
Finally, both protein sequences and token sequences exhibit cross-sequence interaction, where related sequences provide mutual context for understanding the structure of proteins and the semantics of code, respectively. For protein sequences, homologous sequences exhibit similar three-dimensional structures. A well-known example involves hemoglobin and myoglobin, whose protein sequences are homologous and belong to the globin family. They share a three-dimensional structure, known as the globin fold, which consists of eight $\alpha$-helices arranged in a compact, globular configuration. While lexically similar token sequences may differ in variable names, statement ordering, or syntactic constructs, they can still implement the same underlying computational logic and yield identical functional behavior.
3.2. From AlphaFold to AlphaCC: Anology of Framework
We draw an analogy between the frameworks of AlphaFold and AlphaCC to emphasize their connection and distinction.
AlphaFold is a representative end-to-end framework that takes a protein sequence as input and predicts its three-dimensional structure. As illustrated in Figure 2, the overall framework consists of three stages: MSA construction for evolutionary relationships modeling, the Evoformer module for feature extraction, and the structure module for final three-dimensional structure prediction. Since the structure module in AlphaFold serves as a decoder that maps learned features to three-dimensional coordinates—a stage irrelevant to code clone detection—AlphaCC focuses on adapting the first two stages to extract semantic features from token sequences.
In the MSA construction stage, AlphaFold builds a multiple sequence alignment for each protein sequence to compensate for the limited evolutionary information available from a single sequence (Zhang et al., 2024b). AlphaCC adopts a similar design, constructing a multiple sequence alignment of code for each token sequence to enhance semantic understanding. To further enrich token representation, AlphaCC additionally introduces type-specific projections and an interaction mechanism.
Based on MSA, AlphaFold employs Evoformer to capture two crucial types of information: positional information of amino acids within each sequence and interactions among aligned amino acids across the entire MSA (Lupo et al., 2022). Inspired by Evoformer, AlphaCC introduces inner-sequence self-attention and inter-sequence cross-attention to process Code MSAs. These two attention modules capture structural dependencies within individual token sequences, such as control and data flows, and identify co-occurrence patterns across aligned code tokens.
In the final stage, the Structure module from AlphaFold is removed. Instead, to tailor AlphaCC for code clone detection, the token-level similarity between code fragments is calculated through late interaction mechanism.
@@ 4. AlphaCC
Building on cross-domain adaptations from AlphaFold, the framework of AlphaCC enables the detection of semantically equivalent code pairs across syntactic and structural variations, with broad applicability across programming languages. To support this, our method is implemented through a three-stage process, which is described in the following sections.
@@ 4.1. Token Semantic Enhancer
Codebase Preparation. To construct Code MSAs for token sequences, a large-scale and high-quality codebase is required. As our experimental datasets (see Section 5) involve programs written in Java and C, we build separate codebases for each language. Specifically, we collect their top 10,000 GitHub repositories, ranked by the criticality score, which reflects a project’s overall impact and importance (cri, 2024; Dou et al., 2024). Since the target datasets are function-level, we parse all source code from these repositories into individual functions. This results in a total of 68,093,823 functions for Java and 129,219,483 functions for C.
To obtain vector representations of tokens, we employ Word2Vec (Mikolov et al., 2013), a widely adopted and computationally efficient embedding technique. Compared to large language models, Word2Vec offers a lightweight and practical alternative for unsupervised token representation learning. All source code, including that from the experimental datasets and the constructed codebases, is first tokenized using javalang(jav, 2022) for Java and pycparser(pyc, 2024) for C. The resulting tokens are then used to train the Word2Vec model, which generates embedding vectors for individual tokens.
@@
Figure 4. Pipeline of Encoding Code MSA in AlphaCC
Construction of Code MSA. Inspired by AlphaFold’s use of Multiple Sequence Alignment (MSA) to reveal conserved evolutionary patterns in proteins, we introduce multiple sequence alignment of code (Code MSA) to uncover shared structural patterns across syntactically diverse code implementations. As shown in Figure 3, for each input code fragment, we employ an efficient $N$-gram-based similarity search to retrieve four lexically similar token sequences from a candidate group. $N$-gram is adopted for its simplicity and efficiency in capturing recurring token subsequences, a key feature of code clones (Smith and Horwitz, 2009; Martinez-Gil, 2024; Nakagawa et al., 2021). Each sequence is represented using $5$-grams and compared via cosine similarity. To ensure uniformity, all sequences are standardized to a fixed length. For shorter codebase sequences, we incorporate surrounding code as contextual padding. Prior work has demonstrated that such context provides crucial information for understanding the purpose and behavior of a code fragment (Cheng et al., 2022; Zhang et al., 2024a; Bansal et al., 2021). All retrieved token sequences, together with the original sequence, constitute a Code MSA.
@@
"Separator", "Identifier", "Operator", "Keyword",
"Modifier", "DecimalInteger", "BasicType", "String",
"Boolean", "Null", "DecimalFloatingPoint",
"Annotation", "HexInteger", "HexFloatingPoint", "OtherType"
Type-Aware Mechanism for Feature Fusion. To further enhance the model’s structural awareness, we introduce a type-aware mechanism. As summarized in Table 1, following previous research (Dou et al., 2024), all tokens are first categorized into 14 lexical types (e.g., Identifier, Operator, Keyword), along with an OtherType that accounts for rare tokens comprising less than 1% of the dataset. Each type is assigned a unique identifier. Based on the standardized Code MSAs, we extract the corresponding token type IDs and transform them into continuous vectors via an embedding layer. In parallel, the tokens themselves are encoded using a pretrained Word2Vec model. The resulting token and type embeddings are combined through element-wise addition, yielding a unified Code MSA matrix that captures both token semantic and structural information.
@@ 4.2. Codeformer
Inner-Sequence Self-Attention. This component operates row-wise across the Code MSA matrix, where each row represents a single token sequence. It captures intra-procedural logic, including control flow (e.g., if-else branches) and data flow (e.g., variable definitions and usages). Instead of computing separate attention maps for each sequence, Codeformer aggregates them into a shared attention map applied across all sequences. This design leverages the structural homology often present among code fragments (as discussed in Section 3.1), allowing the model to identify consistent interaction patterns that would be overlooked by isolated self-attention.
@@ 4.3. Classification and Optimization
The final stage is responsible for comparing the encoded representations and optimizing the model for the clone detection task.
@@
Table 2. Comparison of different methods on three datasets
GCJ BCB OJClone
Type Method Precision Recall F1 Score Precision Recall F1 Score Precision Recall F1 Score
SourcererCC 74.8 4.2 7.9 97.5 0.6 1.2 27.2 25.2 26.2
NIL 41.0 9.7 15.7 86.4 3.0 5.8 18.0 37.8 24.4
Token-based Toma 68.3 29.1 40.8 80.7 20.7 32.9 71.1 34.0 46.0
PNIAT+PDG 93.4 94.9 94.1 92.0 95.9 93.8 - - -
Graph-based PNIAT+CFG 93.8 96.0 94.9 92.4 96.6 94.5 - - -
Tamer 60.7 10.3 17.5 86.0 2.1 4.1 - - -
TBCCD 97.3 92.7 95.0 96.6 93.1 94.8 95.4 96.8 96.1
ASTNN 97.9 93.0 95.4 97.9 92.8 95.3 97.5 98.3 97.9
AST-based DSFM 97.2 96.3 96.8 95.3 96.0 95.7 95.7 99.4 97.5
Ours AlphaCC 97.6 97.2 97.4 97.7 94.2 96.0 99.7 97.6 98.6
Type Method T1 T2 ST3 MT3 T4
Token-based SourcererCC 100 97.1 63.7 6.2 0
NIL 100 97.4 68.3 20.5 0
Toma 100 100 94.3 87.3 28.1
Graph-based PNIAT+PDG 100 100 99.2 98.9 91.3
PNIAT+CFG 100 100 100 99.2 93.9
AST-based Tamer 100 100 98.3 51.2 2.7
TBCCD 100 100 100 98.3 94.5
ASTNN 100 100 100 99.7 95.1
DSFM 100 100 100 100 95.4
Ours AlphaCC 100 100 100 100 95.7
@@ 5. Evaluation
  • RQ2 (Efficiency): How efficient is AlphaCC in terms of both training and prediction time, particularly for large-scale tasks?
@@
  • RQ4 (Robustness and Generalization): How robust is AlphaCC’s performance when trained and tested on datasets with different characteristics and data distributions?
@@ 5.1. Experiment Setup
Datasets. We conducted experiments on three widely-used datasets. These datasets are strategically chosen to cover different programming languages and diverse clone types, ensuring a thorough and comprehensive evaluation.
  • GCJ (Google Code Jam) (gcj, 2017): A Java dataset derived from the Google Code Jam programming competition, comprising $1,664$ Java projects drawn from $12$ different competition problems. In line with prior work (Zhao and Huang, 2018; Xu et al., 2024), two code snippets that deal with the same problem are treated as clones. It has been confirmed that most clone pairs in GCJ are semantic clone (Zhao and Huang, 2018). We randomly select $1,000$ code snippets for training, $332$ for validation, and $332$ for testing, resulting in nearly $600,000$ code pairs overall.
  • BigCloneBench (big, 2020): A massive Java dataset, which is constructed from IJaDataset (Svajlenko et al., 2014) and contains over $6$ million clone pairs and $260,000$ false clone pairs. All the clone pairs are hierarchically classified into different clone types: Type-1 (T1) clones are lexically identical after comment removal and formatting normalization; Type-2 (T2) clones extend T1 by permitting systematic renaming of identifiers and replacement of literals. The remaining clone pairs are further categorized based on their similarity scores at the line and token levels: Strongly Type-3 (ST3, similarity $\in[0.7,1.0)$) and Moderately Type-3 (MT3, similarity $\in[0.5,0.7)$), while Type-3+4 (T4) exhibit semantic similarity with minimal textual overlap (similarity $\in[0.0,0.5)$(Svajlenko et al., 2014). T4 clones correspond to the notion of semantic clones as emphasized in the GCJ dataset. The scale and diversity of BigCloneBench make it well-suited for evaluating models across both syntactic and semantic clone detection challenges. Following prior work (Wei and Li, 2017; Yu et al., 2019; Xu et al., 2024), we randomly select $8,134$ code snippets for training, $500$ for validation, and $500$ for testing. Finally, we get $124,750$ code pairs for validation and test set and $750,500$ for training set, including $0.3\%$ of T1 clones, $0.1\%$ of T2 clones, $0.2\%$ of ST3 clones, $1.0\%$ of MT3 clones, and $98.4\%$ of T4 clones.
@@
Baselines. To comprehensively evaluate the effectiveness of AlphaCC, we select some state-of-the-art baselines across three types of approaches: token-based, AST-based, and graph-based.
@@
Experimental Settings. In our experiments, all baseline methods are configured using default settings, widely adopted hyperparameters, or values tuned on our datasets to ensure fair comparison. For our model, we set the number of training epochs to $1$, the learning rate to $0.0001$, the margin parameter to $0.5$, and the token embedding dimension to $256$. All experiments are conducted on a unified computational platform equipped with a 192-core CPU, 755 GiB of RAM, and an NVIDIA L40S GPU to ensure consistency and fairness in comparison.
@@ 5.2. RQ1: Effectiveness Comparison
It can be seen that AlphaCC consistently achieves the highest F1 scores across all three distinct benchmarks, reaching $97.4\%$ on GCJ, $96.0\%$ on BigCloneBench, and an outstanding $98.6\%$ on OJClone. This consistent superiority across datasets with diverse programming languages and clone types highlights the effectiveness of our approach. Moreover, AlphaCC maintains top-tier precision and recall, demonstrating its balanced ability to both detect true clones and avoid false positives.
Semantic Clone Detection Superiority. The primary challenge in modern clone detection is accurately identifying semantic clones, which exhibit functional equivalence despite significant syntactic divergence. On two semantic clone datasets, GCJ and OJClone, AlphaCC demonstrates outstanding performance. To further analyze AlphaCC’s strength, we conduct a fine-grained analysis on the challenging BigCloneBench dataset, which categorizes clone types by their syntactic similarity (Table 3). AlphaCC still demonstrates leading performance across all clone types. Crucially, for the most difficult T4 clones, AlphaCC achieves a remarkable F1 score of 95.7%, significantly surpassing all other methods. This performance gain stems from two core architectural innovations: Code MSA, which enhances semantic representation by aligning lexically similar token sequences to reveal functional commonalities; and the Codeformer module, which jointly captures inner-sequence semantics and inter-sequence commonalities, enabling robust abstraction of underlying program logic.
Inherent Multi-Language Applicability. Furthermore, AlphaCC’s core design offers a crucial advantage: inherent multi-language applicability. As shown in Table 2, methods dependent on language-specific analyzers, such as PNIAT, fail to process the C-based OJClone dataset. In contrast, AlphaCC operates on universal token sequences, converting source code into a standardized token stream that abstracts away language-specific syntax. This language-agnostic approach enables AlphaCC to deliver outstanding performance on OJClone, highlighting its potential for deployment across diverse software ecosystems.

### v2

@@ 3.2. Encoder Adaptation
In protein sequences, the role of an amino acid depends on its neighboring residues. For instance, in the lysozyme protein sequence, the segment from positions $25$ to $35$ forms an $\alpha$-helix structure and consists of the following amino acids:
@@
where every letter denotes an amino acid and the accompanying number specifies its position. The relative positioning of amino acids such as W28 and W31 contributes to the formation of a hydrophobic core that stabilizes the $\alpha$-helix structure, while the local arrangement of W28–V29–A30 creates a hydrophobic patch that is essential for local stability.
Similarly, in a token sequence, the semantic role of each token is determined by its surrounding token context. For example, considering the C‑style loop
@@
Based on this property, AlphaFold designs Evoformer to capture two crucial types of information: positional relationships of amino acids within each sequence and interactions among aligned amino acids across the entire MSA (Lupo et al., 2022). Inspired by Evoformer, AlphaCC introduces Codeformer, an encoder designed to model interactions among tokens both within individual token sequences and across aligned sequences in the Code MSA. Codeformer adopts a dual-attention architecture analogous to Evoformer: inner-sequence self-attention captures structural dependencies within individual token sequences, such as control and data flows, while inter-sequence cross-attention identifies co-occurrence patterns across aligned code tokens.
@@ 4. AlphaCC
Building on cross-domain adaptations from AlphaFold, the framework of AlphaCC enables the detection of semantically equivalent code pairs across syntactic and structural variations. To support this, our method is implemented through a three-stage process, which is described in the following sections.
@@ 4.1. Token Semantic Enhancer
Codebase Preparation. To construct Code MSAs for token sequences, a large-scale codebase is required. As our experimental datasets (see Section 5) involve programs written in Java and C, we build separate codebases for each language. Specifically, we collect their top 10,000 GitHub repositories, ranked by the criticality score, which reflects a project’s overall impact and importance (7; S. Dou, Y. Wu, H. Jia, Y. Zhou, Y. Liu, and Y. Liu (2024)). Since the target datasets are function-level, we parse all source code from these repositories into individual functions. This results in a total of 68,093,823 functions for Java and 129,219,483 functions for C.
@@
Construction of Code MSA. Inspired by AlphaFold’s use of multiple sequence alignment (MSA) to reveal conserved evolutionary patterns in proteins, we introduce multiple sequence alignment of code (Code MSA) to uncover shared structural patterns across syntactically diverse code implementations. As shown in Figure 3, for each input code fragment, we employ an efficient $N$-gram-based similarity search to retrieve lexically similar token sequences. $N$-gram is adopted for its simplicity and efficiency in capturing recurring token subsequences, a key feature of code clones (Smith and Horwitz, 2009; Martinez-Gil, 2024; Nakagawa et al., 2021). It serves as a lightweight and scalable retrieval method to ensure efficient candidate alignment and broad contextual coverage. Following previous methods (Nakagawa et al., 2021), each sequence is represented using $5$-grams and compared via cosine similarity to retrieve the top-$k$ most similar sequences. To ensure uniformity, all sequences are standardized to a fixed length. All retrieved token sequences, together with the original sequence, constitute a Code MSA.
@@
"Separator", "Identifier", "Operator", "Keyword", "Modifier", "DecimalInteger", "BasicType", "String", "Boolean", "Null", "DecimalFloatingPoint", "Annotation", "HexInteger", "HexFloatingPoint", "OtherType"
Type-Aware Mechanism for Feature Fusion. To further enhance the model’s structural awareness, we introduce a type-aware mechanism. As summarized in Table 1, following previous research (Dou et al., 2024), all tokens are first categorized into 14 lexical types (e.g., Identifier, Operator, Keyword), along with an OtherType that accounts for rare tokens comprising less than 1% of the dataset. Each type is assigned a unique identifier. Based on the standardized Code MSAs, we extract the corresponding token type IDs and transform them into continuous vectors via an embedding layer. The resulting token and type embeddings are combined through element-wise addition, yielding a unified Code MSA matrix that captures both token semantic and structural information.
@@ 4.2. Codeformer
Inner-Sequence Self-Attention. This component operates row-wise across the Code MSA matrix, where each row represents a single token sequence. It captures intra-procedural logic, including control flow (e.g., if-else branches) and data flow (e.g., variable definitions and usages). Instead of computing separate attention maps for each sequence, Codeformer aggregates them into a shared attention map applied across all sequences. This design leverages the similar functional patterns often present among cloned code fragments, allowing the model to identify consistent interaction patterns that would be overlooked by isolated self-attention.
@@ 4.3. Classification and Optimization
The final stage is responsible for comparing the encoded representations and optimizing the model for the code clone detection task.
@@ 5. Evaluation
  • RQ2 (Efficiency): How efficient is AlphaCC, particularly for large-scale tasks?
@@
  • RQ4 (Robustness and Generalization): How well does AlphaCC generalize across different data distributions and cover instances unsupported by existing methods?
@@ 5.1. Experiment Setup
Datasets. We conducted experiments on three widely used datasets. These datasets are strategically chosen to cover different programming languages and diverse clone types, ensuring a thorough and comprehensive evaluation.
  • GCJ (Google Code Jam) (14): A Java dataset derived from the Google Code Jam programming competition, comprising $1,664$ Java projects drawn from $12$ different competition problems. In line with prior work (Zhao and Huang, 2018; Xu et al., 2024), two code snippets that deal with the same problem are treated as clones. It has been confirmed that most clone pairs in GCJ are semantic clone (Zhao and Huang, 2018). We randomly select $1,000$ code snippets for training, $332$ for validation, and $332$ for testing, resulting in nearly $600,000$ code pairs overall.
  • BigCloneBench (5): A massive Java dataset, which is constructed from IJaDataset (Svajlenko et al., 2014) and contains over $6$ million clone pairs and $260,000$ false clone pairs. All the clone pairs are hierarchically classified into different clone types: Type-1 (T1) clones are lexically identical after comment removal and formatting normalization; Type-2 (T2) clones extend T1 by permitting systematic renaming of identifiers and replacement of literals. The remaining clone pairs are further categorized based on their similarity scores at the line and token levels: Strongly Type-3 (ST3, similarity $\in[0.7,1.0)$) and Moderately Type-3 (MT3, similarity $\in[0.5,0.7)$), while Type-3+4 (T4) exhibit semantic similarity with minimal textual overlap (similarity $\in[0.0,0.5)$(Svajlenko et al., 2014). T4 clones correspond to the notion of semantic clones as emphasized in the GCJ dataset. The scale and diversity of BigCloneBench make it well-suited for evaluating models across both syntactic and semantic clone detection challenges. Following prior work (Wei and Li, 2017; Yu et al., 2019; Xu et al., 2024), we randomly select $8,134$ code snippets for training, $500$ for validation, and $500$ for testing. Finally, we get $124,750$ code pairs for validation and test set and $750,500$ for training set, including $0.3\%$ of T1 clones, $0.1\%$ of T2 clones, $0.2\%$ of ST3 clones, $1.0\%$ of MT3 clones, and $98.4\%$ of T4 clones.
@@
Baselines. To comprehensively evaluate the effectiveness of AlphaCC, we select some state-of-the-art baselines across two types of approaches: token-based and IR-based. For IR-based methods, we mainly select from AST-based and graph-based approaches.
@@
Experimental Settings. In our experiments, all baseline methods are configured using default settings, widely adopted hyperparameters, or values tuned on our datasets to ensure fair comparison. For our model, we set the number of training epochs to $1$, the learning rate to $0.0001$, the margin parameter to $0.5$, the token embedding dimension to $256$, and the top-$k$ value in Code MSA to $4$. All experiments are conducted on a unified computational platform equipped with a 192-core CPU, 755 GiB of RAM, and an NVIDIA L40S GPU to ensure consistency and fairness in comparison.
@@ 5.2. RQ1: Effectiveness Comparison
Table 2. Comparison of different methods on three datasets
GCJ BCB OJClone Type Method Precision Recall F1 Score Precision Recall F1 Score Precision Recall F1 Score SourcererCC 74.8 4.2 7.9 97.5 0.6 1.2 27.2 25.2 26.2 NIL 41.0 9.7 15.7 86.4 3.0 5.8 18.0 37.8 24.4 Token-based Toma 68.3 29.1 40.8 80.7 20.7 32.9 71.1 34.0 46.0 PNIAT+PDG 93.4 94.9 94.1 92.0 95.9 93.8 - - - IR (Graph)-based PNIAT+CFG 93.8 96.0 94.9 92.4 96.6 94.5 - - - Tamer 60.7 10.3 17.5 86.0 2.1 4.1 - - - TBCCD 97.3 92.7 95.0 96.6 93.1 94.8 95.4 96.8 96.1 ASTNN 97.9 93.0 95.4 97.9 92.8 95.3 97.5 98.3 97.9 IR (AST)-based DSFM 97.2 96.3 96.8 95.3 96.0 95.7 95.7 99.4 97.5 Ours AlphaCC 97.6 97.2 97.4 97.7 94.2 96.0 99.7 97.6 98.6
AlphaCC demonstrates superior performance across diverse benchmarks, confirming its effectiveness in code clone detection by achieving the highest F1 scores on all three datasets: $97.4\%$ on GCJ, $96.0\%$ on BigCloneBench, and an outstanding $98.6\%$ on OJClone. This consistent superiority across datasets with diverse programming languages and clone types highlights the effectiveness of our approach. Moreover, AlphaCC maintains top-tier precision and recall, demonstrating its balanced ability to both detect true clones and avoid false positives. Critically, to the best of our knowledge, AlphaCC represents the first token-based method reported to demonstrate superior performance over IR-based approaches in semantic clone detection.
To further explore the performance of AlphaCC, we conduct a fine-grained analysis on the challenging BigCloneBench dataset, which categorizes clone types by their syntactic similarity (Table 3). AlphaCC still demonstrates leading performance across all clone types. Specifically, AlphaCC achieves a perfect F1 score of $100\%$ on T1, T2, ST3, and MT3 clones, matching the highest scores obtained by other methods. Crucially, for the most difficult T4 semantic clones, which involve significant syntactic variations, AlphaCC achieves a remarkable F1 score of 95.7%. This result is the highest among all evaluated methods, notably surpassing strong IR-based baselines. This exceptional performance gain stems from two core architectural innovations: Code MSA, which enhances semantic representation by aligning lexically similar token sequences to reveal functional commonalities; and the Codeformer module, which jointly captures inner-sequence semantics and inter-sequence commonalities, enabling robust abstraction of underlying program logic and effectively addressing the challenge of identifying semantic clones.
Type Method T1 T2 ST3 MT3 T4
Token-based SourcererCC 100 97.1 63.7 6.2 0
NIL 100 97.4 68.3 20.5 0
Toma 100 100 94.3 87.3 28.1
IR (Graph)-based PNIAT+PDG 100 100 99.2 98.9 91.3
PNIAT+CFG 100 100 100 99.2 93.9
IR (AST)-based Tamer 100 100 98.3 51.2 2.7
TBCCD 100 100 100 98.3 94.5
ASTNN 100 100 100 99.7 95.1
DSFM 100 100 100 100 95.4
Ours AlphaCC 100 100 100 100 95.7
5.3. RQ2: Efficiency Analysis
To investigate RQ2, we evaluate AlphaCC’s computational efficiency, a key requirement for practical deployment on large-scale clone detection tasks. We benchmark all methods on one million randomly selected code pairs from BigCloneBench, reporting the average execution time over three runs. Following prior work (Yu et al., 2023; Dou et al., 2024), efficiency comparisons are concentrated on three stages: data preprocessing stage, training stage (required by deep learning-based methods), and prediction stage. The evaluation results are reported in Table 4.